Supermarket CRM system has past records;
a supermarket finds that 26% of people who enter the supermarket will make a purchase. 18 people enter the supermarket during a one-hour period.
(a) What is the probability that exactly 10 customers, 18 customers and 3 customers make a purchase?
(b) Find the expected number of customers who make a purchase.
(c) Find the variance and standard deviation of the number of customers who make a purchase.
1
Expert's answer
2020-08-03T19:16:26-0400

Since the experiment consists of n identical trials.Each trial results in one of the two outcomes, called success and failure.The probability of success, denoted p, remains the same from trial to trial.The n trials are independent ,we use binomial distribution.

Formula: P(x)= "nC_{x}p^{x}q^{n-x}"


a)P(10)= "nC_{10}" "p^{10}q^{n-10}" = "18C_{10}*0.26^{10} *0.74^{18-10} \u2248 0.00555"


P(18)= "nC_{18}" "p^{18}q^{n-18}" ="18C_{18}*" "0.26^{18} *0.74^{18-18}" ≈ 0.0000000000295.


P(3)= "nC_{3}" "p^{3}q^{n-3}" = "18C_{3}*" "0.26^{3}* 0.74^{18-3}" ≈ 0.157


b)Expected value µ = np = 18 * 0.26 = 4.68


c) "Variance: \u03c3^\n2 = npq = 18 *0.26 * 0.74 = 3.46"

"Standard Deviation: σ =\sqrt{npq}= \sqrt{3.46}


 = 1.86"